OpenStudy (anonymous):

x2 + 6x -55 = 0 and root1 = 5, root2 = -11

mathslover (mathslover):

\[x2 + 6x -55 = 0\] \[x^2+11x-5x-55=0\] \[x(x+11)-5(x+11)=0\] \[(x-5)(x+11)=0\] \[x=5 , x=-11\] correct

OpenStudy (anonymous):

then what does a= b= c=

mathslover (mathslover):

where is a , b and c in this equation ?

OpenStudy (anonymous):

oh sorry x2 + 6x -55 = 0 and root1 = 5, root2 = -11 a = , b = , c =

mathslover (mathslover):

sorry i didn;t get u where is a b and c in this eqn ?

OpenStudy (anonymous):

thats what they are asking for a,b,c

mathslover (mathslover):

i think either question is missing or it is wrong

OpenStudy (anonymous):

nope I can snip the equation

OpenStudy (anonymous):

thats the full equation

mathslover (mathslover):

a = 1 , b = 1 , c = -55

mathslover (mathslover):

is it showing it as right ??

OpenStudy (anonymous):

its partial credit

mathslover (mathslover):

i can not understand by partial credit ?

OpenStudy (asnaseer):

the general form of a quadratic equation is:\[ax^2+bx+c=0\]your equation is given as:\[x^2 + 6x -55 = 0\]so compare the coefficients to work out a, b and c.

OpenStudy (anonymous):

some of its right, but something is wrong But I figured it out the answer is a = 1 , b = 6 , c = -55

OpenStudy (asnaseer):

yes - that is the correct answer aussy123.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

both of you!!

OpenStudy (asnaseer):

yw :)

mathslover (mathslover):

oops big mistake sorry :(

OpenStudy (anonymous):

no prob. im just happy you helped

mathslover (mathslover):

:) thanks this is your gr8ness :)

OpenStudy (anonymous):